when a tree 7.2m tall casts a shadow 4.8m long, how long is a shadow cast by a person 1.5m tall?
Answer by stanbon(75887)

when a tree 7.2m tall casts a shadow 4.8m long, how long is a shadow cast by a person 1.5m tall?
------
shadow/object = shadow/object
---
x/1.5 = 4.8/7.2
----
x = (1.5*4.8)/7.2
----
x = 1 meter
====================
Cheers,
Stan H.